OpenStudy (anonymous):

a line passes through the point (3,6) and has slope m=0. find the equation in point-slope form. A)y-3=0 B)y+6=0 C)y+3=0 D)y-6=0

hero (hero):

Hint: y = mx + b (6) = (0)(3) + b 6 = b